two different cards are chosen at random from a standard deck of 52 cards find the probability that both cars are hearts given that both cards are the same color
Answer #1

probability that both cars are hearts given that both cards are the same color =P(both are hearts)/P(both are same color)

=(13/52)*(12/51)/(25/51)=3/25